Two rainstorms occurred in one week in a certain area. The first storm lasted 30 hours, and the second storm lasted 20 hours, for a total 1700mL of rain. What was the rate of rainfall for each of the two storms if the sum of the two rates was 65mL per hour?

Answer by ankor@dixie-net.com(22740)

Two rainstorms occurred in one week in a certain area. The first storm lasted 30 hours, and the second storm lasted 20 hours, for a total 1700mL of rain. What was the rate of rainfall for each of the two storms if the sum of the two rates was 65mL per hour?
:
Let x = rainfall rate of the 1st storm
Let y = rainfall rate of the 2nd
:
Total rainfall equation
30x + 20y = 1700
:
Rain sum per hr equation
x + y = 65
or
y = (65-x)
;
Substitute above for y in the 1st equation:
30x + 20(65-x) = 1700
30x + 1300 - 20x = 1700
30x - 20x = 1700 - 1300
10x = 400
x = 40 ml/hr is the 1st storm
and
65-40 = 25 ml/hr is the 2nd storm
:
:
Check solutions in the total rainfall equation
30(40) + 20(25) =
1200 + 500 = 1700, confirms our solutions